Under a certain condition , cane sugar in water is converted into dextrose at a rate proportional to the amount that is unconverted at any time . If 75 kg at t =; 0, 8 kg are converted during the first 30 minutes , find the amount converted in 2 hours .

1
Expert's answer
2022-03-07T17:26:05-0500

Let "m_0 = 75kg" be the initial mass of the shugar, "m" - is the amount of converted sugar. Thus, the differential equation of the process will be:


"\dfrac{dm}{dt} = r(m_0-m)"

where "r" is the coefficient of proportionality. Solving the euqiation, obtain:


"m = m_0(1 - e^{-rt})"

Since "m = 8kg" at "t = 0.5hrs", obtain the coeffitient "r":


"r = -\dfrac{1}{t}\ln(1-\dfrac{m}{m_0} ) = -\dfrac{1}{0.5hrs}\ln(1-\dfrac{8kg}{75kg} ) \approx 0.23hrs^{-1}"

At "t = 2hrs" obtain:


"m = 75kg\cdot (1-e^{-0.23hrs^{-1}\cdot 2hrs})\approx 48kg"

Answer. 48kg.